Grade 8General Physics

A pure inductor of 25.0 mH is connected to a source of 220 V. Find the inductive reactance and rms current in the circuit if the frequency of the source is 50 Hz

prasanjeet kumar
12 Years agoGrade 8

1 Answer

Saurabh Koranglekar
6 Years ago
The following simple steps will lead you to resolve this query:

1) The resistance of the inductive coil is its inductive reactance which is given by the formula XL= 2πfL.

2) Substitute the values of f and L to get XL. You should get the answer as 7.85 Ω

3) Calculate peak current through the inductor using the formula I0= V/XL= 28 A

4) Now, finally calculate thermscurrent as Irms= I0/√2 = 19.8 A